Someone tries to sell you a solid but irregularly
Answer #1

8) Upthrust = Reduction is weight x g = (25.95 - (71.36-50))g=4.59g=44.982 N

9) Volume of solid = Upthrust/ Density of Water = 44.982/1000g=0.00459 m^3

   density = m/V=25.95/0.00459=5653.59 kg/m^3

10) .Avg Density = Density of Gold x (Total Volume - Hollow Volume)/Total Volume (Neglecting the density of hollow space)

Hollow Volume= (1-(Avg Density/Density of Gold))x Total VOlume= 0.00324 m^3